11. (Section 7.9 and Chapter 8) Consider the 3–dimensional vector field F defined by F (x, y, z) = ￾ 12x 2 y 2 + 2z 2 + 1, 8x 3 y − 3z, 4xz − 3y − 3  (a) Write down the Jacobian matrix JF (x, y, z). (2) (b) Determine the divergence div F(x, y, z). (2) (c) Determine curl F(x, y, z). (2) (d) Give reasons why F has a potential function. (Refer to the relevant definitions and theorems in the study guide.) (2) (e) Find a potential function of F, using the method of Example 7.9.1. Note, however that that example concerns a 2-dimensional vector field, so you will have to adapt the method to be suitable for a 3-dimensional vector field. Pay special attention to the notation that you use for derivatives of functions of more than one variable.


1
Expert's answer
2021-05-07T09:31:59-0400

Given field,

"F(x,y,z)=(12x^2y^2+2z^2+1,8x^3y-3z,4xz-3y-3)"


or "\vec{F}=(12x^2y^2+2z^2+1)\hat{i}+(8x^3y-3z)\hat{j}+(4xz-3y-3)\hat{k}"


(1) Jacobian Matrix "J(F(x,y,z)=\begin{bmatrix}
 \dfrac{df_x}{dx} & \dfrac{df_x}{dy}& \dfrac{df_x}{dz} \\\\
 \dfrac{df_y}{dx}& \dfrac{df_y}{dy}& \dfrac{df_y}{dz}\\\\
 \dfrac{df_z}{dx} & \dfrac{df_z}{dy}& \dfrac{df_z}{dz}
\end{bmatrix}"



="\begin{bmatrix}
 24xy^2& 24x^2y& 4z \\\\
 24xy^2 & 8x^3& -3\\\\
4z& -3&0
\end{bmatrix}"


(2) "div F(x, y, z)"


"= (\dfrac{d}{dx}\hat{i}+\dfrac{d}{dy}\hat{j}+\dfrac{d}{dz}\hat{k}).((12x^2y^2+2z^2+1)\hat{i}+(8x^3y-3z)\hat{j}+(4xz-3y-3)\hat{k})"


"=24xy^2+8x^3+4x"


(3) "curl F(x,y,z)=\begin{bmatrix}
 \hat{i} & \hat{j} &\hat{k}\\\\
 \dfrac{d}{dx} &\dfrac{d}{dy}&\dfrac{d}{dz} \\\\
 12x^2y^2+2z^2+1& 8x^3y-3z& 4xz-3y-3
\end{bmatrix}"


"=\hat{i}(-3-(-3))-\hat{j}(4z-4z)+\hat{k}(24x^2y-24x^2y)\\

 =0"


(4) F is a potential function because, Whenever There exist a field, There exist a potential associated

with it. We can calculate the potential function F by the integraing the given vector field.


(5) As we know-

Potential function is given by-

"F_v=-\int F.dr"


"=\int ((12x^2y^2+2z^2+1)\hat{i}+(8x^3y-3z)\hat{j}+(4xz-3y-3)\hat{k})(dx\hat{i}+dy\hat{j}+dz\hat{z})"


"=(4x^3y+2z^2x+x+4x^3y^2-3yz+2x^2z-3xy-3x)"


"=8x^3y^2+2x^2z+2xz^2-3xy-3yz-2x"